OpenStudy (cicilybailey14):

Fan+Medal Is the equation a direct variation? If it is, find the constant of variation. y + 8 = -x No Yes ; -1 Yes ; 8 Yes ; 1/8

OpenStudy (anonymous):

y=-x-8 constant -1

OpenStudy (phi):

direct variation means it looks like this: y = k x y is x times some number (k) if you have other stuff, then it's not.

OpenStudy (cicilybailey14):

So wait the answers A?

OpenStudy (michele_laino):

sorry how is defined a direct variation?

OpenStudy (phi):

y + 8 = -x can be written as y = -x - 8 that extra -8 makes it "not direct variation"

OpenStudy (michele_laino):

ok! then right answer is A.

OpenStudy (cicilybailey14):

lol Yes^